INFRARED INTENSITY AS A SPECTRAL SIGNATURE OF H-BOND FORMATION IN SMALL AMMONIA CLUSTERS STUDIED IN HE DROPLETS

\underline{MIKHAIL N. SLIPCHENKO}, 1605 Gilman Hall, Department of Chemistry, Iowa State University, Ames, IA 50011-3111; BORIS G. SARTAKOV, on leave from: General Physics Institute, Russian Academy of Sciences, Moscow, 119991, Russia; ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089, USA.

Infrared spectra and intensities of the ν_{1}, ν_{3} H-stretching modes of (NH_{3})_{n} clusters (n = 2-4) have been obtained using the helium droplet isolation technique. Formation of a hydrogen bond in ammonia dimers leads to an increase of the infrared intensity by about a factor of four. In the larger clusters the infrared intensity per bond is close to that of dimers and approaches the value in the NH3 crystal. Intensity of the overtone 2ν_{4} band in trimers increases by a factor of 10 relative to that in single molecules and in dimers, and is comparable to the intensity of the ν_{1}, ν_{3} fundamental bands in larger clusters. These results provide new insights into the hydrogen bonding in ammonia clusters.